OpenStudy (anonymous):

if one point on a line is (5,-4) and the line's slope is -7, find the y-intercept.

OpenStudy (johnweldon1993):

\[\large y = mx+b\] slope = m = -7 x = your x-value = 5 y = your y-value = -4 and you need to solve for 'b' (the y-intercept... \[\large -4 = -7(5)+b\] can you solve this?

OpenStudy (anonymous):

You go to the point (5,-4) and you go down 7 to the right 1 ( because the slope is negative 7) You should find where the line crosses the y axis and that's the y intercept. Hopefully that helps :)

OpenStudy (anonymous):

thanks @johnweldon1993 how do i salve the problem -4=-7(5)+b?

OpenStudy (johnweldon1993):

Well we just solve for 'b' \[\large -4 = -7(5) + b\] we know that -7 times 5 is -35 so \[\large -4 = -35 + b\] and finally we want to isolate 'b' right? so add 35 to both sides of the equation \[\large b = ?\]